OpenStudy (anonymous):

The population of town a increases from 1,346 to 1,500. In the same period, the population of town B increases from 1,546 to 1,800.

OpenStudy (anonymous):

A. Find the percent of increse in population for A and for town B. B. Which town had the greater percent of increase?

OpenStudy (compassionate):

a) increase in population = 1500 - 1346 = 154 percentage increase = 154/1346 (100%) = 11.44% Do b) the same way, now.